# Beyond Matryoshka: Revisiting Sparse Coding for Adaptive Representation

### v2

@@ 4 Empirical Analysis
[…]
@@ 4.1 Retrieval Time Comparison with MRL
[…]
@@
[…]
4.2 Effect of Input Embedding Dimension $\mathbb{R}^{d}$
Experiment Setup.
We conduct experiments to investigate the relationship between fidelity, input embedding dimension $\mathbb{R}^{d}$, and sparsity under a fixed hidden dimension $\mathbb{R}^{h}$ across different architectures. For ViT backbones, we select ViT-S/16 $(d=384)$ and ViT-L/16 $(d=1024)$, setting the hidden dimension to $h=4096$. For ResNet backbones, we choose Resnet18 $(d=512)$ and Resnet50 $(d=2048)$, with the hidden dimension set to $h=8192$. Other parameters are set to their default values. A more detailed experiment setup is provided in Section E.1.
Analysis.
As shown is Figure 4, our empirical analysis reveals an intriguing pattern: as the input embedding dimension increases, performance degradation diminishes Our model exhibits reduced performance degradation. This observation aligns closely with the Johnson-Lindenstrauss Lemma (Johnson et al., 1984), which posits that mappings from high-dimensional to high-dimensional spaces more effectively preserve structural integrity. This insight is particularly significant, as larger embedding sizes generally encode richer information, thereby achieving better downstream performance. By leveraging these high-dimensional embeddings, our approach more effectively retains essential features and relationships within the data.
Figure 4: CSR achieves higher fidelity as the embedding dimensions of the pre-trained models increase. Fidelity percentages of various architectures (ViT and ResNet backbones) at different sparsity levels reveal a consistent trend.
@@ 4.3 Effect of Hidden Representation Dimension $\mathbb{R}^{h}$
To explore the effect of the hidden representation dimension $\mathbb{R}^{h}$ on our model, we use ViT-Large and ResNet50 as pre-trained backbones, following the analysis in Section 4.2. We vary the hidden representation dimension $h$ from $d$ to $16d$ while keeping all other parameters at their default values. Additional implementation details are provided in Section E.2.
@@
Figure 5 compares model performance across different hidden dimensions under varying sparsity constraints. Notably, a shift in the performance trend occurs at $h=4d$. When $h<4d$, performance gradually improves with increasing hidden dimension, reaching its peak at $h=4d$. However, beyond this point, further increases in $h$ lead to performance degradation, particularly under higher sparsity constraints. This trend aligns with the observations of (Gao et al., 2024), which suggest that excessively large hidden dimensions may not be fully utilized, ultimately diminishing model performance. A similar pattern is observed in ResNet. Based on these findings, we set $h=4d$ as the default configuration for all subsequent experiments unless otherwise specified.
Figure 5: CSR exhibits a rise-and-fall trend across different models and hidden dimensions. Model performance across varying hidden dimensions shows that performance peaks at $h=4d$ but degrades beyond this, especially with higher sparsity.
Figure 6: Comparison of dead latent fractions across loss combinations under varying sparsity constraints. Results show that even equipped with $\mathcal{L}_{\text{auxk}}$ and Multiple-TopK at extreme sparsity levels (i.e., k=8,16,32). CSR further alleviates this issue, outperforming baselines and demonstrating its robustness. Evaluations span TopK values from 8 to 1024 using ResNet50 on ImageNet.
@@ 4.4 Effect of Different Losses
To investigate the impact of different loss functions on model capacity, particularly in addressing the dead latents problem discussed in Section 3.2.1, we conduct experiments using ResNet50 as the backbone. Following the findings in Section 4.3, we set $h=4d$, while keeping all other parameters at their default values.
Figure 7: (Left & Middle): Results [^3] We use weights from (Kusupati et al., 2022) to measure MRL and MRL-E’s performance.of ImageNet 1-NN accuracy across active dimensions (a) and Normalized Retrieval Time (b). Figure 2 (a) shows that CSR  and SAE outperform all baselines across active dimensions, with CSR  particularly excelling in extreme low-dimensional scenarios (active dim = 8, 16, 32), mitigating performance degradation significantly. Figure 2 (b) highlights CSR ’s superior accuracy-compute trade-off, maintaining pre-trained model performance while achieving high retrieval efficiency through sparse representation. (Right): Task-Specific Evaluation results of CSR-32 and baselines are shown in Figure 2 (c). For each task, the model is trained on three datasets and evaluated on three unseen datasets. The embeddings learned by CSR  effectively preserve the generalization capability of the pre-trained model, demonstrating its robustness and practical applicability across diverse tasks.
@@
[…]
@@ 5 Benchmark Results and Analysis
We evaluated the effectiveness of our proposed CSR framework across three mainstream representation modalities: vision, language, and vision+language. For vision representation (see Section 5.1), we conduct image classification on ImageNet-1K and evaluate performance using 1-NN accuracy, following (Kusupati et al., 2022). For language representation (see Section 5.2), we focus on three primary tasks: text classification, text clustering, and text retrieval on the MTEB benchmark (Muennighoff et al., 2022). For multimodal representation (see Section 5.3), we report Recall5 performance on two widely-used datasets: MS COCO (Lin et al., 2014) and Flickr30K (Young et al., 2014). Through these experiments, we aim to provide a holistic understanding of the capabilities of our proposed framework.
5.1 Vision Representation Comparision
Experiment Setup.
We evaluate 1-NN accuracy on ImageNet1k classification, following (Kusupati et al., 2022). The normalized retrieval time is computed through settings in Section 4.1. We select a pre-trained ResNet50 model from (Wightman, 2019) and trained CSR on its encoded embedding of ImageNet1k. For further implementation details, please refer to Section B.
Table 1: Retrieval Efficiency vs. Performance trade-offs under Matched Conditions. We use NV-Embed-V2 as our pre-trained model, and present its performance in the first line of the table in gray. We analyze Dataset-Specific Evaluation results along two key dimensions: (1) Relative Retrieval Time under matched performance and ii) performance under matched retrieval efficiency. Under matched performance, CSR achieves a remarkable 61× speedup, while under matched retrieval efficiency, it improves performance by 15%, demonstrating its superior balance between speed and accuracy. The maximum values are indicated in bold, while the second-highest values are underlined. Relative Retrieval Time is calculated follows the definition in Section 4.
@@ 5.1 Vision Representation Comparision
Method Dim Time MTOPIntent Banking77 TweetSentiment BiorxivP2P BiorxivS2S TwentyNews FiQA2018 NFCorpus SciFACT
NV-Embed-V2 4096 37.6 93.58 92.20 79.73 53.61 49.60 64.82 62.65 43.97 77.93
Stella-en-1.5B-v5 256 2.6 90.45 86.14 76.75 50.81 46.42 60.07 55.59 36.97 77.48
@@
Gecko-Text-Embedding-004-256 256 2.4 77.82 86.01 72.97 36.28 33.09 50.60 55.54 37.81 70.86
OpenAI-Text-Embedding-3-L-256 256 2.8 70.45 83.19 58.98 35.43 33.86 54.24 50.33 37.94 73.10
Arctic-Embed-L-V2.0 256 2.6 67.69 80.99 59.06 34.25 34.07 30.06 44.69 35.02 69.51
@@
SAE (w/ NV-Embed-V2) 32 1.0 87.43 88.11 75.19 51.02 48.68 58.63 49.18 35.14 66.04
CSR (w/ NV-Embed-V2) 32 1.0 89.86 91.02 78.55 53.49 49.13 63.05 57.54 37.06 71.17
@@
Figure 2(a) and (b) illustrate the comparison of learned representation quality through the 1-NN classification accuracy of ResNet50 models trained and evaluated on ImageNet-1K. Reconstruction-based sparse compression methods (CSR & SAE) outperform MRL and MRL-E due to their training flexibility, enabling application top of powerful pre-trained models. These methods also surpass traditional post-hoc compression techniques (e.g., SVD) and linear probes on random features by increasing the overall model total capacity while keeping active dimensions for each sample unchanged, as discussed in Section 1 and Section 3.2.1. This enhanced capability allows CSR to maintain remarkable robustness, with only 2% performance degradation (73.84 vs 75.16). These results highlight that the proposed CSR design can effectively compress pre-trained embeddings while leveraging the natural benefits of sparse matrix multiplication. More detailed experimental results can be found in Section B.4.
@@ 5.2 Text Representation Comparision
We assessed CSR on three key tasks from the MTEB benchmark (Muennighoff et al., 2022), testing it across six datasets for each task. In detail, we conduct evaluations in two distinct settings: Dataset-Specific Evaluation, where CSR is trained and tested on different splits of the same dataset to ensure consistency, and Task-Specific Evaluation, where CSR is trained on one dataset and evaluated on unseen datasets within the same task to rigorously assess its generalization capabilities. We choose NV-Embed-V2 (Lee et al., 2024a) as our pre-trained model and present its performance in gray. For further experimental details, please refer to Section C. We refer to CSR-K as a model with the TopK activations and so as SAE.
@@
Table 1 demonstrates the performance of CSR and baseline models across multiple tasks and datasets. CSR not only maintains the strong performance of the pre-trained model but also surpasses baselines under varying resource constraints. Notably, it achieves a 61× speedup at matched retrieval efficiency and a 15% performance improvement at matched computational cost, highlighting its superior balance between speed and accuracy. Furthermore, as shown in Figure 2(c), CSR preserves the generalization capabilities of the pre-trained model, ensuring robust performance across diverse downstream tasks. These results underscore the efficacy and versatility of CSR , demonstrating its strong potential for real-world applications.
@@ 5.3 MultiModal Representation Comparision
We evaluate our methods on multimodal retrieval tasks using the widely-used MS COCO and Flickr30K datasets, employing the ViT-B-16 backbone. We conduct a MRL fine-tuning on these datasets as baselines, following the standard MRL training paradigm (Kusupati et al., 2022). The performance of our backbone, using the same fine-tuning procedure, is shown in gray. During training, both SAE and CSR leverage a shared sparse embedding layer for images and text. Additional experimental setup and implementation details are provided in Section D.
@@
Table 2 presents the multimodal retrieval task results across different methods. In general, reconstruction-based methods exhibit relatively low performance degradation on both datasets. Compared to the MRL method, CSR achieves average performance gains of 4.6% and 6.8% on image-to-text retrieval, and 9.1% and 6.5% on text-to-image retrieval across the two datasets. SAE experiences more severe performance degradation compared to CSR, which underlines the efficacy of our design in image-text alignment. However, as the sparsity constraint becomes more stringent, the performance gap between CSR and MRL narrows. Upon further investigation, we find that CSR still suffers from the “dead latents” problem even when equipped with advanced mechanisms. Addressing the mitigation of dead latents in the alignment space remains an open challenge, leaving room for future work and study. For a detailed analysis, please refer to Section D.4.
Table 2: MultiModal Retrieval Recall@5 (%) on image-to-text (I2T) and text-to-image (T2I) retrieval across two benchmark datasets, MS COCO and Flickr30K. CSR outperforms ViT-B-16-MRL across all active dimensions and tasks.
MS COCO Flickr30K
Method Active Dim I2T T2I I2T T2I
ViT-B-16 512 74.42 86.47 91.92 97.79
ViT-B-16-MRL 256 67.12 77.53 80.41 89.89
SAE 71.21 82.58 87.76 95.59
CSR 71.41 83.49 87.98 96.79
ViT-B-16-MRL 128 64.19 73.02 77.56 87.80
SAE 64.67 76.70 81.40 91.20
CSR 69.34 81.04 84.05 93.00
ViT-B-16-MRL 64 62.61 72.43 74.22 84.79
SAE 56.30 69.45 70.58 81.30
CSR 62.75 78.10 76.44 88.50
6 Conclusion

### v4

@@ 4 Empirical Analysis
In this section, we conduct a careful study on the empirical performance of the proposed CSR. All experiments in this section are conducted on ImageNet, using 1-NN accuracy (Johnson et al., 2019) as the evaluation metric. By default, we set the hidden dimension $h$ of CSR to be $h=4d$, where $d$ is the dimension of the pretrained dense embeddings, and set the default active dimension to $k=32$.
For a fair and intuitive comparison of MRL and CSR, First, we adopt the notion of active dimension as a surrogate metric to benchmark the retrieval time under dense (MRL-type) and sparse (CSR-type) embeddings. For example, “$\mathrm{Active\ Dim}=8$” denotes either a length-$8$ dense embedding (MRL) or a sparse embedding with TopK ($k=8$) activation (CSR). Notably, we choose it because dense and sparse matrix multiplication have the same computation complexity under the same active dimension $k$, i.e., $\mathcal{O}(k)$. In Section 4.1, we further carefully benchmark them in practice and find that the two indeed have similar retrieval time, and sparse ones can be even slightly faster under small $k$.
To account for variations in retrieval time due to sample size, we establish a standardized benchmarking protocol (denoted as $\mathcal{T}$) to measure retrieval latency by default. Specifically, to simulate large-scale retrieval scenarios, we report the average retrieval time for 512 queries over an ImageNet-scale database containing 1.3 million entries (equivalent to the size of the ImageNet training set). For CSR, we use a default hidden dimension of $h=16{,}384$ and an active dimension of $k=32$. All experiments are conducted in a consistent GPU environment using PyTorch (Paszke et al., 2019). To facilitate comparison, we also report the relative retrieval time of each method by normalizing it against the retrieval time of CSR under the default setup. Additional implementation details can be found in Section E.3.
@@ 4.1 Retrieval Time Comparison with MRL
In this section, we benchmark the retrieval time of MRL and CSR under the same active dimension $k$ and analyze the impact of hidden dimension $\mathbb{R}^{h}$, database size $N$ and sparsity $k$.
(i) Active dimension. Figure 3(a) shows retrieval time under varying hidden dimensions, with database size fixed. We can see that the retrieval time of CSR (i.e., sparse multiplication) and MRL (i.e., dense multiplication) both grow with large $k$ and remain relatively on the same level. And for smaller $k$, CSR shows a clearer advantage over MRL. Although CSR and MRL have similar theoretical complexity $O(dk)$, their actual runtimes are affected by backend implementations. For instance, cuBLAS (used for dense ops) is highly optimized but has high launch overhead, while cuSPARSE (used for CSR) is lighter but less optimized for small $k$. Interestingly, we can observe that for sparse embeddings, retrieval time decreases as hidden dimension $h$ increases. This suggests notable benefit of CSR that it can use higher latent dimensions for better expressivity while attaining faster retrieval. On the contrary, MRL with higher dense dimensions always has slower retrieval. We elaborate potential reasons on this distinction at Appendix E.4.
(ii) Database size. Figure 3(b) shows that CSR demonstrates superior scalability as the database size $N$ increases from $0.5$M to $10$M. The relative efficiency gain becomes more pronounced with larger datasets, underscoring the practicality of sparse embeddings in real-world retrieval scenarios.
Figure 4: Performance of CSR under different sparsity levels with different sizes of backbone models. CSR achieves higher fidelity at greater sparsity levels when applied to larger backbone models (which provide better base performance), observed consistently in both ViT and ResNet architectures.
4.2 Effect of Backbone Size
@@
We examine fidelity versus backbone size (with different input dimension $\mathbb{R}^{d}$), and sparsity, using fixed hidden dimension $\mathbb{R}^{h}$ across architectures. For ViT, we use ViT-S/16 $(d=384)$ and ViT-L/16 $(d=1024)$ with $h=4096$. For ResNet, we test RN18 $(d=512)$ and RN50 $(d=2048)$ with $h=8192$. A more detailed experiment setup is provided in Section E.1.
@@
Figure 4 demonstrates that a larger backbone with higher input embedding dimensions improves model fidelity at equal sparsity levels. This insight is particularly significant, as larger embedding sizes generally encode richer information, thereby achieving better downstream performance. By leveraging these high-dimensional embeddings, our approach more effectively retains essential features and relationships within the data.
Figure 5: Performance of CSR under different hidden dimensions and different types of backbone models (ResNet-50 (convolution) and ViT-L (Transformers)). CSR exhibits a reverse U-shape across different models and hidden dimensions. CSR’s performance peaks at $h=4d$ ($d$ is the input dimension size) but degrades beyond this, especially with higher sparsity.
Figure 6: Comparison of dead latent fractions across loss combinations under varying sparsity constraints. Results show that even equipped with $\mathcal{L}_{\text{auxk}}$ and Multiple-TopK at extreme sparsity levels (i.e., $k=8,16,32$). CSR further alleviates this issue, outperforming baselines and demonstrating its robustness.
Figure 7: (Left & Middle): Results of ImageNet Top-1 accuracy (a) and 1-NN accuracy (b) across active dimensions under the same pretrained ResNet-50 backbone used in Kusupati et al. (2022). We can see that while MRL trains the whole network and CSR only uses frozen embeddings, CSR still performs consistently better across all embedding sizes and has significant margins beyond 20% at lower active dimensions (the region that yields the largest efficiency gains). (Right): Comparison of text embedding methods at similar retrieval cost. For CSR, we use $k=32$ by default. For each task, the model is trained on three datasets and evaluated on three unseen datasets. The text embeddings learned by CSR outperformed other MRL-based baselines by significant margins across different natural language tasks at much lower training cost.
@@ 4.3 Effect of Hidden Representation Dimension $\mathbb{R}^{h}$
We explore how hidden dimension $\mathbb{R}^{h}$ effects on our model, we use ViT-Large and ResNet50 as pre-trained backbones, sweeping $h$ from $d$ to $16d$ while keeping all other parameters at their default values. Additional implementation details are provided in Section E.2.
@@
Figure 5 compares model performance across different hidden dimensions under varying sparsity constraints. Notably, a shift in the performance trend occurs at $h=4d$. When $h<4d$, performance gradually improves with increasing hidden dimension, reaching its peak at $h=4d$. However, beyond this point, further increases in $h$ lead to performance degradation, particularly under higher sparsity constraints. This trend aligns with the observations of  Gao et al. (2024), which suggest that excessively large hidden dimensions may not be fully utilized, ultimately diminishing model performance. A similar pattern is observed in ResNet. Based on these findings, we set $h=4d$ as the default configuration for all subsequent experiments unless otherwise specified.
@@ 4.4 Effect of Different Losses
We investigate how different loss functions affect model capacity, particularly in addressing the dead latent problem discussed in Section 3.2.1, using RN50 backbone with $h=4d$. Other parameters are set at their default values.
@@
Figure 6 illustrates the impact of different loss functions on model capacity. The naïve SAE suffers from severe dead latents, while the inclusion of an auxiliary loss $\mathcal{L}_{\text{aux}}$ and the multi-TopK loss partially mitigates this issue. Introducing a non-negative contrastive loss (NCL) further alleviates the problem, particularly at extreme sparsity levels (e.g., $k=8,16,32$). Empirical results validate the effectiveness of Theorem 5, demonstrating that representation learning with NCL promotes more orthogonal and disentangled features. This, in turn, increases the number of active dimensions and enhances overall model performance.
@@ 5 Benchmark Results and Analysis
We evaluated the effectiveness of our proposed CSR framework across three mainstream representation modalities: vision, language, and vision+language. For vision representation (see Section 5.1), we conduct image classification on ImageNet-1K and evaluate performance using 1-NN accuracy, following Kusupati et al. (2022). For language representation (see Section 5.2), we focus on three primary tasks: text classification, text clustering, and text retrieval on the MTEB benchmark (Muennighoff et al., 2022). For multimodal representation (see Section 5.3), we report both in-distribution and zero-shot cross-modal retrieval performance on two widely-used datasets: MS COCO (Lin et al., 2014) and Flickr30K (Young et al., 2014). Through these experiments, we aim to provide a holistic understanding of the capabilities of our proposed framework.
Table 1: Performance and efficiency of text embeddings on three natural language tasks: classification, clustering, and retrieval. We use NV-Embed-V2 as our pre-trained model, and present its performance in the first line of the table in gray. We analyze Dataset-Specific Evaluation results along two key dimensions: (1) Relative Retrieval Time under matched performance and ii) performance under matched retrieval efficiency. Under matched performance, CSR achieves a remarkable 61× speedup, while under matched retrieval efficiency, it improves performance by 15%, demonstrating its superior balance between speed and accuracy. The maximum values are indicated in bold, while the second-highest values are underlined. Relative Retrieval Time is calculated follows the definition in Section 4.
@@
Category Model Dim Time MTOPIntent Banking77 TweetSentiment BiorxivP2P BiorxivS2S TwentyNews FiQA2018 NFCorpus SciFACT
Full Rep NV-Embed-V2 4096 37.6 93.58 92.20 79.73 53.61 49.60 64.82 62.65 43.97 77.93
MRL Stella-1.5B-v5 256 2.6 90.45 86.14 76.75 50.81 46.42 60.07 55.59 36.97 77.48
@@
Gecko-Embed-004(Google) 256 2.4 77.82 86.01 72.97 36.28 33.09 50.60 55.54 37.81 70.86
Text-Embed-3-L (OpenAI) 256 2.8 70.45 83.19 58.98 35.43 33.86 54.24 50.33 37.94 73.10
Arctic-Embed-L-V2 256 2.6 67.69 80.99 59.06 34.25 34.07 30.06 44.69 35.02 69.51
@@
Sparse SAE (w/ NV-Embed-V2) 32 1.0 87.43 88.11 75.19 51.02 48.68 58.63 49.18 35.14 66.04
CSR (w/ NV-Embed-V2) 32 1.0 89.86 91.02 78.55 53.49 49.13 63.05 57.54 38.06 71.17
5.1 Vision Representation Comparision
Baselines
We compare our proposed method with the following baseline approaches. 1) MRL/MRL-E (Kusupati et al., 2022): RN50 model where the fully connected layer is replaced by multiple (MRL) or a single (MRL-E) classification head(s) that take truncated input dimensions (e.g., only the first 8 of the original 2048 dimensions). 2) SVD: We performed a low-rank approximation of the 1000-way classification layer of RN50, with rank = 1000. 3) Rand-LP: We compared against a linear classifier fit on randomly selected features (He et al., 2020). 4) Rand-FS: We randomly selected features extracted from RN50 for 1-NN classification.
Experiment Setup.
We evaluate 1-NN accuracy and Top-1 accuracy on ImageNet1k classification, following Kusupati et al. (2022). For fair comparison, we used the same RN50 backbone weights as MRL (denoted as FF2048 in the original work) and trained CSR on its ImageNet1k encoded embeddings. For further implementation details, please refer to Section B.
@@ 5.1 Vision Representation Comparision
Figure 7(a) and (b) illustrate the comparison of learned representation quality through the Top-1 and 1-NN classification accuracy of RN50 models trained and evaluated on ImageNet-1K. For linear probing results (Figure 7(a)), reconstruction-based sparse compression methods (CSR & SAE) outperform MRL-LP (both linear probing methods) by a large margin and also surpass MRL/MRL-E (train from scratch) in lower active dim ($k<128$). Furthermore, Figure 7(a) demonstrates the superior representation quality learned by CSR, which consistently outperforms MRL across various active dimensions. CSR also surpass traditional post-hoc compression techniques (e.g., SVD) and linear probes on random features by increasing the overall model total capacity while keeping active dimensions for each sample unchanged, as discussed in Section 1 and Section 3.2.1. This enhanced capability allows CSR to maintain remarkable robustness, even under extrem sparsity where $k=2,4,8$. These results highlight that the proposed CSR design can effectively compress pre-trained embeddings while leveraging the natural benefits of sparse matrix multiplication. More detailed experimental results can be found in Section 4.
@@ 5.2 Text Representation Comparision
We assessed CSR on three key tasks from the MTEB benchmark, testing it across six datasets for each task. In detail, we conduct evaluations in two distinct settings: Dataset-Specific Evaluation, where CSR is trained and tested on different splits of the same dataset to ensure consistency, and Task-Specific Evaluation, where CSR is trained on one dataset and evaluated on unseen datasets within the same task to rigorously assess its generalization capabilities. We choose NV-Embed-V2 (Lee et al., 2024a) as our pre-trained model and present its performance in gray. For further experimental details, please refer to Section C. To improve readability, we refer to CSR-K as a model with the TopK activations and so as SAE.
Table 2: Comparison of different methods on multi-modal retrieval tasks using two benchmark datasets, MS COCO and Flickr30k, evaluated under both in-distribution and zero-shot settings, with Recall@5 (%) as the performance metric. We use ViT-B/16 as our pre-trained model, and present its performance in the first line of the table in gray. For Zero-Shot setting, CSR is first trained on a large-scale scale, dataset-CC3M and evaluated on downstream tasks. CSR (plug-and-play) consistently outperforms ViT-B-16-MRL (fully fine-tuned) in various tasks with significant training efficiency.
In-Distribution Zero-Shot
Active Trainable MS COCO Flickr30K MS COCO Flickr30K
Method Dim Parms I2T T2I I2T T2I I2T T2I I2T T2I
ViT-B-16 512 86M 74.42 86.47 91.92 97.79 69.23 83.03 89.82 97.70
ViT-B-16-MRL 256 86M 67.12 77.53 80.41 89.89 56.90 65.82 80.94 89.20
SAE 1.1M 71.21 82.58 87.76 95.59 58.22 67.40 82.44 86.19
CSR 1.1M 71.41 83.49 87.98 96.79 61.85 70.14 85.22 91.10
ViT-B-16-MRL 128 86M 64.19 73.02 77.56 87.80 53.63 61.16 77.67 85.10
SAE 1.1M 64.67 76.70 81.40 91.20 53.20 63.02 77.54 85.19
CSR 1.1M 69.34 81.04 84.05 93.00 54.37 68.04 78.08 88.09
ViT-B-16-MRL 64 86M 62.61 72.43 74.22 84.79 47.47 54.42 71.16 79.00
SAE 1.1M 56.30 69.45 70.58 81.30 44.48 53.56 69.58 82.29
CSR 1.1M 62.75 78.10 76.44 88.50 48.61 61.90 73.04 84.10
@@
Table 1 demonstrates the performance of CSR and baseline models across multiple tasks and datasets. CSR not only maintains the strong performance of the pre-trained model but also surpasses baselines under varying resource constraints. Taking text classification as an example, CSR achieves a 15% accuracy improvement at matched computational cost (i.e., with retrieval times comparable to Jina-V3-64 and Nomic-Embed-V1.5-64) while attaining a 61x speedup when matched for performance (i.e., compared to NV-Embed-V2). The results underscore CSR ’s exceptional ability to maintain an optimal speed-accuracy trade-off - a critical requirement for practical deployment in large-scale retrieval systems. We further evaluate the generalization capability of CSR (with $k=32$) on three unseen datasets per task, as shown in Figure 7(c). The results demonstrate that sparse representations yield more robust performance compared to dense alternatives at same activation dimensions. These results underscore the efficacy and versatility of CSR , demonstrating its strong potential for real-world applications.
@@ 5.3 MultiModal Representation Comparision
We evaluated our methods on multimodal retrieval tasks using the ViT-B-16 backbone, testing both in-distribution and zero-shot cross-modal retrieval on MS COCO (Lin et al., 2014) and Flickr30K (Young et al., 2014) datasets. For baselines, we fine-tuned MRL on these datasets (using CC3M (Changpinyo et al., 2021) for zero-shot training), following standard MRL training protocols  (Kusupati et al., 2022). The performance of our backbone, using the same fine-tuning procedure, is shown in gray. During training, both SAE and CSR leverage a shared sparse embedding layer for images and text. Additional experimental setup and implementation details are provided in Section D.
@@
Table 2 presents the multimodal retrieval task results across different methods and settings. In general, reconstruction-based methods exhibit relatively low performance degradation on both datasets. Compared to the MRL method, CSR achieves average performance gains of 4.6% and 6.8% on I2T retrieval, and 10.3% and 6.5% on T2I retrieval across the two datasets in In-Distribution Evaluation. Besides, under zero-shot scenario, CSR also surpasses MRL by 3.2% and 3.3% on I2T, and 9.2% and 3.9% on T2I, respectively. Notably, these results demonstrate CSR’s potential to handle large-scale datasets (e.g., CC3M-3M images, compared to ImageNet’s 1M and MS COCO’s 0.3M), confirming CSR’s consistent superiority across various active dimensions and its scalability. SAE experiences more severe performance degradation compared to CSR, which underlines the efficacy of our design in image-text alignment. However, as the sparsity constraint becomes more stringent, the performance gap between CSR and MRL narrows. Upon further investigation, we find that CSR still suffers from the “dead latents” problem even when equipped with advanced mechanisms. Addressing the mitigation of dead latents in the alignment space remains an open challenge, leaving room for future work and study. For a detailed analysis, please refer to Section D.4.
6 Conclusion & Discussion